# Design of high-strength, radiopure copper-chromium alloys for rare-event searches assisted by computational thermodynamics

### v1

@@ Abstract
Direct Dark Matter detection and studies on the nature of neutrinos demand detector systems with extremely low background levels, including from radioactivity. Additive-free, electroformed copper, in addition to a set of advantages, exhibits exceptional radiopurity, making it the material of choice for rare-event searches. To satisfy the increasing demand for materials with superior mechanical strength, the development of copper-chromium alloys is pursued. Early investigations explored the synthesis of these alloys by electrodeposition and thermal processing. A materials-design approach is proposed to optimize the fabrication and thermal processing stages of manufacturing. It is assisted by materials modeling tools based on the thermodynamic and kinetic properties of alloy compositions, which enables faster development of novel materials by predicting properties and materials performance. This approach is demonstrated by comparing simulations with previously reported experimental investigations and proposing improved thermal processing.
@@ 1 Introduction
Immense progress has been achieved in understanding the Cosmos over more than 45 orders of magnitude in length. However, fundamental questions remain about the nature of dark matter (DM) and neutrinos. The existence of DM, corresponding to 84.5% of matter in the Universe, is established via astrophysical observations and precise cosmological measurements [1]. Several search approaches are used, with “direct detection” being one of the most prominent: detect DM from the Milky Way halo through its coherent elastic scattering off a nucleus [2]. At the same time, clarification of whether neutrinos are Majorana or Dirac particles will have profound implications and may lead to the discovery of new physics phenomena. A most sensitive probe is the search for neutrinoless double $\beta$-decay, a process only possible for Majorana neutrinos [3].
Copper is the material of choice for detector systems used for rare-event searches. It is commercially available at low cost and high purity, and has no long-lived radioactive isotopes. However, it may be contaminated during manufacturing and can be activated by fast neutrons from cosmic rays.
@@
To meet the unique radiopurity requirements, attention is focused on electroformed copper (EFCu), which has favorable radiochemical, thermal, and electrical properties. Rare-event search experiments pioneer large-scale, additive-free Cu electroformation, e.g. the Majorana demonstrator [4] and the ECuME project [5]. This novel technique leads to extremely low contamination below 10-14 grams of 232Th and 238U per Cu gram [4, 6].
@@
Despite its advantages, Cu is highly ductile and of low strength, which limits its use for moving mechanical, high-pressure, and load-bearing parts. Alloying Cu with other elements, e.g. Be, Cr, Zr, Ti, Ni, Si can enhance its mechanical properties [7, 8, 9]. This is of interest for a broad range of applications, including transportation, electronics and electric circuits [8, 10, 11, 12]. However, during manufacturing the alloy needs to undergo a thermal strengthening process, which may affect its electrical properties. Thus, a trade-off between strength and electrical conductivity needs to be made, in addition to the balance between strength and radiopurity, relevant for rare event searches.
@@
Since Cr can be plated on the Cu surface and visa versa, copper-chromium (CuCr) alloys have the potential to fulfill the enhanced strength and radiopurity requirements, and are of particular interest for rare-event searches. For example, the DarkSPHERE experiment [13, 14], proposes to use a large-scale high-pressure fully electroformed underground spherical proportional counter to probe uncharted territory in the search for sub-GeV particle DM. An alloy of enhanced strength would enable detector operation at higher pressure and improved physics reach. Another example is the nEXO experiment, a neutrinoless-double $\beta$-decay experiment aiming at clarifying the nature of neutrinos [15], which requires a pressure vessel capable of containing liquid xenon. This could be realised through a high-strength, radiopure alloy compared to the currently proposed titanium construction.
@@
We aspire to use materials modeling tools, based on computational thermodynamics, to design the manufacturing of CuCr alloys, including electroformation and thermal processing, with optimal parameters. In this work, as a first step, simulations are compared with results of published experimental investigations, and improved processes are suggested. To the author’s knowledge this is the first time computational thermodynamics is applied in the design of materials for fundamental particle physics research.
@@ 2 Experimental investigations of CuCr alloys
Cr electroplating takes place at a higher voltage compared to Cu and poses challenges for the material’s radiopurity [16, 6, 17]. The advantage of Cu plating at much lower voltage, i.e. approximately 0.3 V, supresses deposits of Th or U, which are present in the electrolytic bath in the form of ions. Cr on the other hand, needs to be plated at higher voltage, e.g. in the vicinity of 0.5 V, with the risk of increasing the Th and U contamination.
Nevertheless, small additions of Cr to Cu lead to substantial enhancement of the CuCr alloy strength [9]. This has also been demonstrated in initial experimental investigations aiming at manufacturing electroformed CuCr alloys [16, 17], while maintaining high radiopurity. Local measurements of CuCr at 0.585 wtCr indicated that it can improve the hardness of EFCu by 70% [16, 18]. A single point measurement of CuCr material with maximum composition of approximately 0.4 wtCr points to maximum hardness above 140 HV; i.e. 100% harder than EFCu found in the same study [17].
@@
Solution heat treatment is required so that a homogenized composition can be achieved in a single phase, before quenching - i.e. cooling at a very fast rate. Pure Cr and Cu are organised thermodynamically at different crystal structures. Cr is found in the body centered cubic (bcc) and Cu in the faced centered cubic (fcc) crystal structure. At a temperature high enough for a single phase to be thermodynamically stable, after phase transformation occurs, Cu and Cr atoms are organized in an fcc crystal structure. Cr atoms diffuse from the bcc structure until the whole volume of the bcc phase is transformed into a single fcc matrix phase, leading to a supersaturated solid solution after quenching. The misfit between the different sizes of the Cu and Cr atoms occupying the sites of the fcc lattice leads to lattice distortion and matrix strengthening by hindering dislocations movement. Besides the difference in the size of atoms of solute in the matrix, the concentration of solute atoms is also important [7, 8].
@@
Aging aims at precipitation strengthening and follows solid solution strengthening. Starting from a fully homogenized alloy composition and a single-phase microstructure, the controlled growth of a secondary phase, i.e. bcc, precipitates into the Cu-rich matrix during aging. The two-phase microstructure is thermodynamically stable in a temperature range lower than that at which homogenization can be achieved. Cr diffuses from the matrix to form spherical precipitates. The incoherence between the two phases and the size of the spherical precipitates are the main reasons for mechanical strengthening. During precipitation strengthening, the formation of precipitates results in a decrease in the concentrations of solute atoms n the matrix, which reduces the impact on electrical conductivity [8].
@@
Aging is an evolving process, where the secondary phase starts as an embryo and gradually reaches the size of the spherical precipitates at a critical radius, at which the size is ideally very small, and the volume fraction of the precipitates is the maximum possible for the specific alloy composition. Solid solution strengthening, previously achieved through solution heat treatment for alloy homogenization, is not as pronounced at this stage, where heat treatment is taking place at a much lower temperature than the homogenization stage. Because Cr atoms will diffuse from the matrix to the precipitating phase, the fcc lattice distortion decreases and so will the yield strength due to solid solution strengthening. By extending aging for longer periods, larger precipitates evolve by simultaneously dissolving smaller precipitates at the coarsening stage, leading to a lower yield strength increase due to precipitation strengthening.
@@
In the process of improving mechanical properties of CuCr alloys for low background application, two strengthening mechanisms are crucial, namely solid solution strengthening and precipitation strengthening. An example from industrial application, demonstrated that Cu-0.5Cr (in wt%) alloy exhibits hardness above 160 HV, tensile strength of 499 MPa, and electrical conductivity of 74.8 %IACS, post aging for 1 hour at 450[9]. In this investigation the alloy was fabricated from casting, solution heat treated for 3 hours at 1050C, and mechanically strengthened through cold rolling. Although casting and mechanical deformation compromise radiopurity, and thus cannot be used in the application at hand, the modeling aspects for thermal processing, especially for the aging stage, are relevant for this investigation.
@@
In order to optimize the parameters of temperature and time for both solution heat treatment and aging stages, a thermodynamic and kinetic description of the Cu-Cr system is required. For example, the solubility of Cr in Cu is limited and only a narrow single-phase field of less than 1 wtcan be found, in the range of 1020C – 1080C pointing to temperatures that could be selected for the solution heat treatment stage [19, 20, 21]. After quenching, the combination of temperature and duration of time needs to be carefully selected to optimize the precipitation strengthening stage with aging. Since the size of precipitates and their distribution in the matrix determine the mechanical enhancement of Cu-based alloys [8, 9, 22], the optimized parameters in the design of the thermal process for specific alloy compositions can be determined using computational thermodynamics [22, 23, 24, 25, 26, 27].
@@ 3 Modeling Methods
Modeling using computational thermodynamics aims to determine the optimal time and temperature for the homogenization and aging stages. These tools are constructed within the CALPHAD framework (CALculation of PHAse Diagrams) [28, 29, 30]. The CALPHAD-type databases Thermo-calc Software TCHEA6 and MOBHEA3 High Entropy alloys database [31, 32] are used in this work for the description of the thermodynamic and kinetic properties of Cu-Cr system.
@@ 3.1 Solution heat treatment (homogenization)
For the solution heat treatment stage, aiming at homogenization, DICTRA 1D simulations are employed [33]. These solve the diffusion equations at the interface of two regions of different composition during phase transformation. This is done at an isothermal, i.e. a single temperature for a specific duration of time, in a volume-fixed frame of reference [34, 35].
The setup consists of two regions, separated by a planar boundary, with initial concentrations of Cr and Cu and phases of bcc and fcc, respectively, at each region. For each of them, the single-phase model [33] is used. Because Cr and Cu are found in different phases, the moving boundary model [33] is applied, where phase transformation occurs at the interface of the two regions. Assuming a local two-phase equilibrium and mass balance at the interface, the simultaneous solution of the equations yields the interfacial concentrations and the growth rates of each species. At the end of each simulation, the concentration profiles of each species are obtained.
The expression of the flux $J_{i}$ of each component (e.g. Cr) can be expanded in terms of concentration gradients $\vartheta{c_{j}}$ in direction $z$:
@@
Finally, the flux-balance equation for mass conservation of each component determines the migration rate at the phase interface:
@@
where $\nu^{bcc/fcc}$ is the interface velocity, $c^{bcc}_{Cr}$ and $c^{fcc}_{Cr}$ are the concentrations of Cr in bcc and fcc phase, respectively, and $J^{bcc}_{Cr}$ and $J^{fcc}_{Cr}$ the fluxes of Cr at each side of the interface. Integration in time includes initial calculation of boundary conditions at the interface, by combining the moving boundary and single-phase models.
@@ 3.2 Aging (precipitation strengthening)
Nucleation and growth of precipitates during aging are simulated using TC-PRISMA [22], according to the Langer-Schwartz theory [36]. In addition, the Kampmann Wagner numerical method is adopted [37]. In this way, the nucleation and growth of precipitating secondary phase are considered simultaneously. At the end of each simulation the mean radius, the volume fraction, and the particle size distribution of precipitates are obtained.
According to the classical nucleation theory, the nucleation rate $J_{s}$ for a steady state is given by:
@@
where $\Delta G^{*}$ is the nucleation barrier, $Z$ is the Zeldovich factor, $\beta^{*}$ the atomic attachment rate, and $N$ the number of nucleation sites. The nucleation barrier is given by:
@@
where $\Delta G_{m}$ is the driving force of the multicomponent system, $V_{m}$ is the molar volume of the precipitating phase and $\sigma$ is the interfacial energy, where the value of 0.7 $J/m^{2}$ is adopted for Cu-0.5Cr (in wt%) following Ref. [9].
@@ 3.3 Solution and precipitation strengthening
where $\chi_{\rm Cr}$ is the mole fraction of the solute species (Cr, in at%) in the matrix, $G$ is the shear modulus of the matrix (for Cu), and $\delta$ and $\eta$ are factors associated with the lattice misfit and the variation due to shear modulus, respectively.
@@
where $b$ the Burgers vector for Cu matrix, $\nu$ the Poisson’s ratio for Cu, $r$ the mean particle radius, $f$ the precipitate volume fraction, and $\bf{\lambda}$ the effective interparticle spacing [9]:
@@
Table 1 lists the inputs used to determine yield strengthening.

### v2

@@ Abstract
Direct Dark Matter detection and studies on the nature of neutrinos demand detector systems with extremely low background levels, including from radioactivity. Additive-free, electroformed copper, in addition to a set of advantages, exhibits exceptional radiopurity, making it the material of choice for mechanical and structural components for rare-event searches experiments. To satisfy the increasing demand for materials with superior mechanical strength, the development of copper-chromium alloys is pursued. Early investigations explored the synthesis of these alloys by electrodeposition and thermal processing. A materials-design approach is proposed to optimize the fabrication and thermal processing stages of manufacturing. It is assisted by materials modeling tools based on the thermodynamic and kinetic properties of alloy compositions, which enables faster development of novel materials by predicting properties and materials performance. This approach is demonstrated by comparing simulations with previously reported experimental investigations and proposing improved thermal processing.
@@ 1 Introduction
Immense progress has been achieved in understanding the Cosmos over more than 45 orders of magnitude in length. However, fundamental questions remain about the nature of dark matter (DM) and neutrinos. The existence of DM, corresponding to 84.5% of matter in the Universe, is established via astrophysical observations and precise cosmological measurements [1]. Several search approaches are used, with “direct detection” being one of the most prominent: to detect DM from the Milky Way halo through its coherent elastic scattering off a nucleus [2]. At the same time, clarification of whether neutrinos are Majorana or Dirac particles will have profound implications and may lead to the discovery of new physics phenomena. A most sensitive probe is the search for neutrinoless double $\beta$-decay, a process only possible for Majorana neutrinos [3].
@@
Copper is the material of choice for structural detector components used for rare-event searches. It is commercially available at low cost and high purity, has excellent thermal and electrical conductivity, and has no long-lived radioactive isotopes. However, it may be contaminated during manufacturing and can be activated by fast neutrons from cosmic rays. To meet the unique radiopurity requirements, attention is focused on electroformed copper (EFCu), which can be plated at an electric potential that suppresses depositions of Th and U contaminants. This novel technique leads to extremely low contamination below 10-14 grams of 232Th and 238U per Cu gram [4, 5]. Rare-event search experiments pioneer large-scale, additive-free Cu electroformation, e.g. the completed Majorana demonstrator [4] and the on-going ECuME project [6]. In addition, to support the construction of experiments, underground laboratories provide manufacturing of EFCu as a service [7].
@@
Despite its advantages, Cu is highly ductile and of low strength, which limits its use for moving mechanical, high-pressure, and load-bearing parts. Alloying Cu with other elements, e.g. Be, Cr, Zr, Ti, Ni, Si can enhance its mechanical properties [8, 9, 10]. This is of interest for a wide range of applications, including transportation, electronics, and electric circuits [9, 11, 12, 13]. However, during manufacturing, the alloy needs to undergo a thermal strengthening process which may affect its electrical properties. Thus, a trade-off between strength and electrical conductivity needs to be made, in addition to the balance between strength and radiopurity, relevant for rare-event searches.
@@
Since Cr can be plated on the Cu surface and vice versa, copper-chromium (CuCr) alloys have the potential to fulfill the enhanced strength and radiopurity requirements, and are of particular interest for rare-event searches. For example, the DarkSPHERE experiment [14, 15], proposes to use a large-scale high-pressure fully electroformed underground spherical proportional counter to probe uncharted territory in the search for sub-GeV particle DM. An alloy of enhanced strength would enable detector operation at higher pressure and improved physics reach. Another example is the nEXO experiment, a neutrinoless-double $\beta$-decay experiment aiming at clarifying the nature of neutrinos [16], which requires a pressure vessel capable of containing liquid xenon. This could also be realized through a high-strength, radiopure alloy.
@@
We aim to use materials modeling tools, based on computational thermodynamics, to design the manufacturing of CuCr alloys, including electroformation and thermal processing, with optimal parameters. In this work, as a first step, simulations are compared with results of published experimental investigations, and improved processes are suggested. To the best of author’s knowledge, this is the first time computational thermodynamics has been applied in the design of materials for fundamental particle physics research. Further work towards synthesis of application specific Cu-based alloys and their impact to rare-event search experiments is reported in Ref. [17].
@@ 2 Experimental investigations of CuCr alloys
The advantage of Cu plating at much lower voltage, i.e. approximately 0.3 V, is that it suppresses deposits of Th or U, which are present in the electrolytic bath in the form of ions. Cr, on the other hand, needs to be plated at higher voltage, e.g., in the vicinity of 0.5 V, with the risk of increasing Th and U contamination. However, CuCr alloys are projected to have radiopurity at levels similar to those of EFCu [18].
Small additions of Cr to Cu lead to a substantial increase in the strength of the CuCr alloy [10]. This has also been demonstrated in initial experimental investigations aiming at manufacturing electroformed CuCr alloys [18, 19], while maintaining a high radiopurity. Local measurements of CuCr at 0.585 wt% Cr indicated that it can improve the hardness of EFCu by 70% [18, 20]. A single point measurement of CuCr material with a maximum composition of approximately 0.4 wt% Cr points to maximum hardness above 140 HV (Vickers hardness); i.e. 100% harder than EFCu found in the same study [19].
@@
Solution heat treatment is required so that a homogenized composition can be achieved in a single phase, before quenching - i.e. cooling at a very fast rate. Pure Cr and Cu are organized thermodynamically in different crystal structures. Cr is found in the body centred cubic (bcc) and Cu in the faced centred cubic (fcc) crystal structure. At a temperature high enough for a single phase to be thermodynamically stable, after phase transformation occurs, Cu and Cr atoms are organized in an fcc crystal structure. Cr atoms diffuse from the bcc structure until the whole volume of the bcc phase is transformed into a single fcc matrix phase, leading to a supersaturated solid solution after quenching. At this point, the solution heat treatment has been completed. The misfit between the different sizes of the Cu and Cr atoms occupying the sites of the fcc lattice leads to lattice distortion and matrix strengthening by hindering dislocations movement. In addition to the difference in the size of atoms of the solute in the matrix, the concentration of the solute atoms is also important [8, 9].
@@
Aging aims at precipitation strengthening and follows solid solution strengthening. Starting from a fully homogenized alloy composition and a single-phase microstructure, the controlled growth of a secondary phase, i.e. bcc, precipitates into the Cu-rich matrix during aging. The two-phase microstructure is thermodynamically stable in a temperature range lower than that at which homogenization can be achieved. Cr diffuses from the matrix to form spherical precipitates. The incoherence between the two phases and the size of the spherical precipitates are the main reasons for mechanical strengthening. During precipitation strengthening, the formation of precipitates results in a decrease in the concentrations of solute atoms in the matrix, which reduces the impact on electrical conductivity [9].
@@
Aging is an evolving process, where the secondary phase starts as an embryo and gradually reaches the size of the spherical precipitates at a critical radius, at which the size is ideally very small, and the volume fraction of the precipitates is the maximum possible for the specific alloy composition. Solid solution strengthening, previously achieved through solution heat treatment for alloy homogenization, is not as pronounced at this stage, where heat treatment is taking place at a much lower temperature than the homogenization stage. Because Cr atoms will diffuse from the matrix to the precipitating phase, the fcc lattice distortion decreases and so will the yield strength due to solid solution strengthening. By extending aging for longer periods, larger precipitates evolve by a simultaneous dissolution of smaller precipitates at the coarsening stage, leading to a lower yield strength increase due to precipitation strengthening.
@@
In the process of improving mechanical properties of CuCr alloys for low background application, two strengthening mechanisms are crucial, namely solid solution strengthening and precipitation strengthening. An example from industrial application, demonstrated that Cu-0.5Cr (in wt%) alloy exhibits hardness above 160 HV, tensile strength of 499 MPa, and electrical conductivity of 74.8 %IACS (International Annealed Copper Standard), after aging for 1 hour at 450[10]. In this investigation, the alloy was fabricated from casting, solution heat treated for 3 hours at 1050C, and mechanically strengthened by cold rolling. Although casting and mechanical deformation compromise radiopurity, and thus cannot be used in the application at hand, the modeling aspects for thermal processing, especially for the aging stage, are relevant for this investigation.
@@
In order to optimize the parameters of temperature and time for both solution heat treatment and aging stages, a thermodynamic and kinetic description of the Cu-Cr system is required. For example, the solubility of Cr in Cu is limited and only a narrow single-phase field of less than 1 wt% can be found, in the range of 1020C – 1080C pointing to temperatures that could be selected for the solution heat treatment stage [21, 22, 23]. After quenching, the combination of temperature and duration of time needs to be carefully selected to optimize the precipitation strengthening stage with aging. Since the size of precipitates and their distribution in the matrix determine the mechanical enhancement of Cu-based alloys [9, 10, 24], the optimized parameters in the design of the thermal process for specific alloy compositions can be determined using computational thermodynamics [24, 25, 26, 27, 28, 29].
@@ 3 Modeling Methods
The modeling using computational thermodynamics aims to determine the optimal time and temperature for the homogenization and aging stages. These tools are constructed within the CALPHAD framework (CALculation of PHAse Diagrams) [30, 31, 32]. DICTRA and TC-PRISMA, presented in the following, are kinetic simulation modules, in the Thermo-calc Software package. In this work, Thermo-Calc version 2024a is used, and the CALPHAD-type databases Thermo-calc Software TCHEA6 and MOBHEA3 High Entropy alloys database [33, 34], for the description of the thermodynamic and kinetic properties of the Cu-Cr system.
@@ 3.1 Solution heat treatment (homogenization)
For the solution heat treatment stage, to achieve homogenization, DICTRA 1D simulations are employed [35]. These solve the diffusion equations at the interface of two regions of different composition during phase transformation. This is done at an isothermal, i.e. a single temperature for a specific duration of time, in a volume-fixed frame of reference [36, 37].
The setup consists of two regions, separated by a planar boundary, with initial concentrations of Cr and Cu and phases of bcc and fcc, respectively, at each region. For each of them, the single-phase model [35] is used. Because Cr and Cu are found in different phases, the moving boundary model [35] is applied, where the phase transformation occurs at the interface of the two regions. Assuming a local two-phase equilibrium and mass conservation at the interface, the simultaneous solution of the equations yields the interfacial concentrations and the growth rates of each species. At the end of each simulation, the concentration profiles of each species are obtained.
In the diffusion equations, the expression of the flux $J_{i}$ of each component (e.g. Cr) can be expanded in terms of concentration gradients $\vartheta{c_{j}}$ in direction $z$:
@@
The flux-balance equation ensuring mass conservation of each component determines the migration rate at the phase interface:
@@
where $\nu^{bcc/fcc}$ is the interface velocity, $c^{bcc}_{Cr}$ and $c^{fcc}_{Cr}$ are the Cr concentrations in the bcc and fcc phases, respectively, and $J^{bcc}_{Cr}$ and $J^{fcc}_{Cr}$ the fluxes of Cr on each side of the interface. Integration in time includes initial calculation of boundary conditions at the interface, by combining the moving boundary and single-phase models.
@@ 3.2 Aging (precipitation strengthening)
The nucleation and growth of precipitates during aging are simulated using TC-PRISMA [24], according to the Langer-Schwartz theory [38]. In addition, the Kampmann Wagner numerical method is adopted [39]. In this way, the nucleation and growth of the precipitating secondary phase are considered simultaneously. At the end of each simulation, the mean radius, the volume fraction, and the particle size distribution of the precipitates are obtained.
According to the classical nucleation theory, on which the Langer-Schwartz theory is based, the nucleation rate $J_{s}$ for a steady state is given by the following:
@@
where $\Delta G^{*}$ is the nucleation barrier, i.e. the energy cost associated with forming a stable nucleus of the new phase, $Z$ is the Zeldovich factor, $\beta^{*}$ the atomic attachment rate, and $N$ the number of nucleation sites. The nucleation barrier is given by:
@@
where $\Delta G_{m}$ is the driving force of the multicomponent system, i.e. the thermodynamic potential that favors the formation of the new phase, $V_{m}$ is the molar volume of the precipitating phase, and $\sigma$ is the interfacial energy, where the value of 0.7 $J/m^{2}$ is adopted for Cu-0.5Cr (in wt%) following Ref. [10].
@@ 3.3 Solution and precipitation strengthening
where $\chi_{\rm Cr}$ is the mole fraction of the solute species (Cr, in at%) in the matrix, $G$ is the shear modulus of the matrix (for Cu), and $\delta$ and $\eta$ are factors associated with the lattice misfit and the variation due to the shear modulus, respectively.
@@
where $b$ the Burgers vector for the Cu matrix, $\nu$ the Poisson’s ratio for Cu, $r$ the mean particle radius, $f$ the precipitate volume fraction, and $\bf{\lambda}$ the effective interparticle spacing [10]:
@@
Table 1 lists the inputs used to determine the yield strengthening.